OpenStudy (anonymous):

PLEASE HELP Solve the given equation! 8^2x=11

OpenStudy (anonymous):

x=11/64

OpenStudy (anonymous):

Okay so, since neither the base nor the exponent is same, we take log log8^2x = log11 By rule of log, 2x.log8=log11 Or, 2x=log11/log8 So, 2x=1.153 x=1.153/2 x=0.576

OpenStudy (anonymous):

i would use this fact \[b^x=A\iff x=\frac{ln(A)}{\ln(b)}\] to write \[2x=\frac{\ln(11)}{\ln(2)}\] then divide by 2